OpenStudy (ohohaye):

How do I solve for t in 0=−5t^2+30t+10

OpenStudy (johnweldon1993):

We want to know when a function equals 0, sounds like we need to know when said function crosses the x-axis. This function is a quadratic in the form \(\large ax^2 + bx + c = 0\) and therefor the quadratic formula can be used to find the roots "zeros" of the function \[\large \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] Plug in your coefficients into this equation and find the solutions to find your 't' values that make this function = 0

OpenStudy (ohohaye):

Oh ok, thank you